# On the Robustness of Adversarial Training Against Uncertainty Attacks

### v1

@@ 1 Introduction
The role of UQ techniques assumes a highly pivotal role in security-sensitive applications: in such scenarios, in fact (e.g., medical diagnostics and autonomous driving [3, 2]), the tolerance on the error margin must be, per definition, minimized and strictly confined. Leveraging UQ techniques, however, ML-based systems can also base their output on the uncertainty estimated on the given input, helping minimize errors and enabling an accurate and reliable decision-making process throughout the ML systems’ pipelines.
@@
Within the same security-related contexts, however, offspring of an instead now mature line of research, ML models are also often required to be resistant against adversarial inputs, i.e., carefully crafted input samples aiming to deceive the classification performed by the model [8, 9]. Following an arms race in the security of ML, the literature has flourished, proposing multiple defense approaches and attack algorithms [10]. Among the State-Of-the-Art (SoA) defense approaches, Adversarial Training (AT) [11] has shown to be the reference technique by also being the pillar of more recent approaches, and its effectiveness is deeply and continuously studied against the SoA adversarial attacks [12].
In contrast, for attacks against uncertainty, little to no work has advanced the study of defensive approaches rather than focusing on different ways to disrupt the UQ techniques. We thus questioned, starting from the principle of evaluating existing approaches first, the effectiveness of AT against attacks to uncertainty. In our work, we reveal, through theoretical analysis and empirical investigation, that models trained to be robust against adversarial attacks can be likewise robust against attacks to uncertainty; we summarize the main theoretical findings of our analysis in  Fig. 1.
Figure 1: An illustration of the effect of standard (left) and adversarial (right) training on the Uncertainty Span (i.e., the range of uncertainty values one can obtain by perturbing a sample with a given budget $\epsilon$) for a classification problem visualized through a standard 3-dimensional simplex. While for standard training, the predicted probability Refer to caption converges on the ground truth probability Refer to caption, for adversarial training, it results in less confidence, i.e., has higher entropy than the ground truth probability. As a result, adversarial training reduces the uncertainty span, whose lower and upper bounds are, respectively, an over-confidence Refer to caption and under-confidence Refer to caption attack. For further details, see  Sect. 3 and  Sect. 4.
In Sect. 2, we display the background concepts of uncertainty quantification and adversarial ML; then, in Sect. 3, starting from the state of the art, we exemplify a modular framework for dealing with adversarial attacks targeting uncertainty, which is functional to the theoretical analysis on the robustness of adversarial trained models of Sect. 4. Successively, in Sect. 5, we empirically demonstrate our results by considering, on multiple datasets, highly robust SoA models from the publicly available RobustBench repository. Finally, in Sect. 6, we outline our findings and discuss the potential limitations of the study at hand.
2 Background
2.1 Uncertainty Quantification
@@
where each $\mathbf{B}_{s}\in\mathcal{D}$ contains all the samples having confidence (i.e.,, $\arg\max f(\boldsymbol{x})$) within the interval $[\frac{s}{S},\frac{s+1}{S}]$, and ${ex}(\mathbf{B}_{s})=\frac{2s+1}{2S}$ is the average over the interval. For measuring epistemic uncertainty, a common approach employs the variance of the predictions: $\mathbb{V}(f^{BMA}(\boldsymbol{x}))\coloneqq\mathbb{E}[(f^{BMA}(\boldsymbol{x}))^{2}]-\mathbb{E}[(f^{BMA}(\boldsymbol{x}))]^{2}$.
2.2 Adversarial Attacks
In a security-sensitive application, another priority, together with reliable uncertainty estimates, is to guarantee a certain level of robustness in a model that can be the target of attackers aiming at disrupting its normal functioning. This necessity brought to light what is known as adversarial machine learning, an ML field that keeps exploring the attack surfaces of an ML system to propose new countermeasures to achieve robustness [10, 14]. In this regard, most of the research community efforts focused on attacks jeopardizing the integrity of a model, where an attacker is interested in changing a model’s prediction at test time [9, 8]. In such an attack scenario, the goal of the attacker is to craft a small perturbation $\boldsymbol{\delta}^{\prime}$ such that, once added to an input sample $\boldsymbol{x}$, it induces the target model $f^{\boldsymbol{\theta}}$ to output a wrong prediction, i.e., $f^{\boldsymbol{\theta}}(\boldsymbol{x}+\boldsymbol{\delta}^{\prime})\neq f^{\boldsymbol{\theta}}(\boldsymbol{x})$. Here, the perturbation is typically constrained to a limited domain $\mathcal{B}_{\epsilon}=\{\forall\boldsymbol{\delta}\ :||\boldsymbol{\delta}||_{p}\leq\epsilon\}$, where ||.$||.||_{p}$ is an arbitrary $L_{p}$ norm (e.g.,, $L_{1}$, $L_{2}$ or $L_{\infty}$) and $\epsilon$ is the perturbation budget indicating the limits of the attacker’s power.
@@ 2.2 Adversarial Attacks
Defending against adversarial attacks – Currently, the most promising approach to protect a model against adversarial attacks is to proactively anticipate the attacker by including them during training. Such procedure, which takes the name of adversarial training (AT) [11], aims at finding the optimal parametrization $\boldsymbol{\theta}^{\prime}$ that ideally protects from all possible perturbations $\boldsymbol{\delta}\in\mathcal{B}_{\epsilon}$. This strategy can be formalized as the following minimax problem:
@@
where $\mathcal{L}$ is the objective that minimize classification error (e.g.,, the cross-entropy loss). Here, Eq. 6 is composed of (i) an inner maximization problem to find, for each sample, the perturbation $\boldsymbol{\delta}$ that achieves high loss, and (ii) an outer minimization problem to find the parameters $\boldsymbol{\theta}$ that minimize this worst-case loss.
3 Adversarial Attacks Against Uncertainty Quantification
From the perspective of the most typical test-time evasion attacks, adversarial machine learning characterizes an attacker interested in making the model misclassify the prediction on the input adversarial example. In the context of Uncertainty Quantification, instead, the focus shifts from “deteriorating predictions” to “deteriorating the uncertainty measure”. The correct functioning of an uncertainty measure implies that the most likely correct predictions have a low uncertainty, which, in contrast, becomes high when the predictions are plausibly incorrect. Therefore, while predictions and uncertainty represent two different measures, they are, however, affine: the uncertainty estimate measures the likelihood that the given prediction is wrong. As we will see in the following subsection, state-of-the-art work explicitly conceived uncertainty adversarial attacks to spoil this statistical correlation.
3.1 Previous Work
@@
In our previous work [6], we proposed an alternative scenario in which the attacker is interested in increasing or decreasing the predictive uncertainty of all the input samples. The ability to damage predictive uncertainty without having specific information about the labels is an attractive feature, especially if such a measure is the input of a downstream module or is given to a human. For instance, imagine a system employing the uncertainty measure of an ML model for determining if the clinical data of a patient is sufficient to make a diagnosis: without knowledge of this specific domain (thus without the labels), an attacker may craft a perturbation for increase indiscriminately the uncertainty associated to the patient’s data. Such an attack may overload the processing pipeline of a hospital, causing the system to require more analysis, even when unnecessary. The proposed attack scenario takes the name of over-confidence (when decreasing the uncertainty) and under-confidence attack (when increasing the uncertainty). While for the over-confidence attack, we kept the same optimization process of  Eq. 8, the under-confidence case has been investigated only theoretically, without an actual loss function implementation.
@@
Obadinma et al.[7] propose to harm the model’s calibration, disrupting the functioning of any downstream module or human using the uncertainty measure. For doing so, they present four different attacks, two of which coincide with the formulation we proposed in  [6], i.e., (i) over-confidence and (ii) under-confidence attack. They further propose (iii) maximum miscalibration attack (MMA), where they increase or decrease the confidence of each sample following the same criteria of  [4], and (iv) random confidence attack (RCA), where they extend the latter but with a randomized confidence goal for each input, producing a more natural confidence distribution that, although being a less effective attack, it results to be more stealthy. Based on these types of attacks against model’s calibration, they propose two ad-hoc defenses: (i) Calibration Attack Adversarial Training (CAAT), in which they perform adversarial training using samples generated with maximum miscalibration attacks, and (ii) Compression Scaling (CS), a post-processing technique where they heuristically distribute low scores to higher range. To the best of our knowledge, this is the only work that proposes a defense strategy against uncertainty attacks.
3.2 A Modular Unified Framework for Uncertainty Attacks
@@
Interestingly, all the state-of-the-art works build an optimization process with a proper loss formulation for accomplishing an attacker’s goal. This means that proving the model’s robustness against a specific goal does not directly translate the same robustness guarantees on other goals. For example, analyzing the effect of an attack on OOD uncertainty minimization does not provide any information on the effectiveness against indiscriminate uncertainty maximization, making it difficult to compare models in terms of uncertainty robustness. Accordingly, we propose a modular approach for unifying the diverse attacks proposed in the literature, which, as a result, also unifies the notion of robustness. In general, all the possible instances of the attacker’s goal have a shared characteristic: they all need to modify the uncertainty measure associated with the model’s predictions; more sensitive uncertainty variations generally lead to less robustness against uncertainty attacks. As a result, we divert our attention not to any specific instance of the attacker’s goal but rather to estimate the range of uncertainty values an attacker can reach when perturbing each input $x$ on the dataset $\mathcal{D}$. Concretely, given an attacker capable of perturbing a sample $\boldsymbol{x}$ inside a domain $\mathcal{B}_{\epsilon}$, we can distinguish between two different cases: (i) decreasing the uncertainty, (i.e., over-confidence attack) and (ii) increasing the uncertainty (i.e., under-confidence attack). These objectives can be achieved, respectively, by minimising or maximising the uncertainty measure $\mathcal{U}(f^{\boldsymbol{\theta}}(\boldsymbol{x}))$:
@@
$\underbrace{\boldsymbol{\delta}^{\downarrow}=\operatornamewithlimits{\arg\,\max}_{\boldsymbol{\delta}\in\mathcal{B}_{\epsilon}}\;\mathcal{U}(f^{\boldsymbol{\theta}}(\boldsymbol{x}+\boldsymbol{\delta}))}_{\textbf{Over-confidence Attack}}\quad\quad\quad\underbrace{\boldsymbol{\delta}^{\uparrow}=\operatornamewithlimits{\arg\,\min}_{\boldsymbol{\delta}\in\mathcal{B}_{\epsilon}}\;\mathcal{U}(f^{\boldsymbol{\theta}}(\boldsymbol{x}+\boldsymbol{\delta}))}_{\textbf{Under-confidence Attack}}$ (9)
@@
In accordance with the state of the art, we use the entropy of the prediction vector $H(f^{\boldsymbol{\theta}}(\boldsymbol{x}))$ [1] as our uncertainty measure $\mathcal{U}(\cdot)$. To instantiate an over-confidence attack, the objective is to minimize the entropy of the output probability. To do so, it is sufficient to force the model to output $f^{\boldsymbol{\theta}}_{k}(\boldsymbol{x}+\boldsymbol{\delta}^{\downarrow})=1$ for an arbitrary $k\in\mathcal{Y}$. However, it is crucial to mention that we are not proposing a new, strong attack implementation; hence, the objective of our work is not to propose a formulation that maximizes the Uncertainty Span. Instead, we focus on analyzing the robustness of ML models against such attacks in general, the reason for which we will use an attack formulation that is theoretically founded and empirically adequate while still not saying that it would be the most effective one. We know from [6] that the most straightforward strategy for the attacker is to maximize the probability of the class corresponding to the predicted class, as it already has the highest probability among all the classes, as in  Eq. 8; this attack takes the name of Stabilizing Attack (STAB) since it stabilizes the model’s prediction by enforcing the most likely class [6]. To achieve this objective, one can employ the cross-entropy loss, setting as target $\boldsymbol{t}^{\downarrow}$ the one-hot-encoding of the predicted label $\hat{y}$:
@@
4 Theoretical Analysis on Adversarial Uncertainty Robustness
Despite much literature on the attacker side, only one work dealt with adversarial robustness against uncertainty attacks, proposing two ad hoc strategies. In contrast to them, driven by the flourishing literature on defenses against prediction attacks (i.e., traditional adversarial examples), before thinking of ad hoc strategies, we ask ourselves if and to what extent such defenses - particularly adversarial training - may protect even against uncertainty attacks. Accordingly, our main research question is:

### v2

@@ I Introduction
The role of UQ techniques assumes a highly pivotal role in security-sensitive applications: in such scenarios, in fact (e.g., medical diagnostics [3, 4] and autonomous driving [2]), the tolerance on the error margin must be, by definition, minimized and strictly confined. Leveraging UQ techniques, however, ML-based systems can also base their output on the uncertainty estimated on the given input, helping minimize errors and enabling an accurate and reliable decision-making process throughout the ML systems’ pipelines.
@@
Within the same security-related contexts, however, offspring of an instead now mature line of research, ML models are also often required to be resistant against adversarial inputs, i.e., carefully crafted input samples aiming to deceive the classification performed by the model [9, 10]. Following an arms race in the security of ML, the literature has flourished, proposing multiple defense approaches and attack algorithms [11]. Among the State-Of-the-Art (SoA) defense approaches, Adversarial Training (AT) [12] has been shown to be the reference technique by also being the pillar of more recent approaches, and its effectiveness is deeply and continuously studied against the SoA adversarial attacks [13].
In contrast, for attacks against uncertainty, little to no work has advanced the study of defensive approaches, rather than focusing on different ways to disrupt the UQ techniques. We thus questioned, starting from the principle of evaluating existing approaches first, the effectiveness of AT against attacks on uncertainty. In our work, we reveal, through theoretical analysis and empirical investigation, that models trained to be robust against adversarial attacks can be likewise robust against attacks to uncertainty; we summarize the main theoretical findings of our analysis in  Fig. 1.
Figure 1: An illustration of the effect of standard (left) and adversarial (right) training on the Uncertainty Span (i.e., the range of uncertainty values one can obtain by perturbing a sample with a given budget $\epsilon$) for a classification problem visualized through a standard 3-dimensional simplex. While for standard training, the predicted probability Refer to caption converges on the ground truth probability Refer to caption, for adversarial training, it results in less confidence, i.e., has higher entropy than the ground truth probability. As a result, adversarial training reduces the uncertainty span, whose lower and upper bounds are, respectively, an over-confidence Refer to caption and under-confidence Refer to caption attack. For further details, see  Sect. III and  Sect. IV.
In Sect. II, we display the background concepts of uncertainty quantification and adversarial ML; then, in Sect. III, starting from the state of the art, we exemplify a modular framework for dealing with adversarial attacks targeting uncertainty, which is functional to the theoretical analysis on the robustness of adversarial trained models of Sect. IV. Successively, in Sect. V, we empirically demonstrate our results by considering, on multiple datasets, highly robust SoA models from the publicly available RobustBench repository. Finally, in Sect. VI, we outline our findings and discuss the potential limitations of the study at hand.
II Background
II-A Uncertainty Quantification
@@
where each $\mathbf{B}_{s}\in\mathcal{D}$ contains all the samples having confidence (i.e., $\arg\max f(\boldsymbol{x})$) within the interval $[\frac{s}{S},\frac{s+1}{S}]$, and ${ex}(\mathbf{B}_{s})=\frac{2s+1}{2S}$ is the average over the interval. For measuring epistemic uncertainty, a common approach employs the variance of the predictions: $\mathbb{V}(f^{BMA}(\boldsymbol{x}))\coloneqq\mathbb{E}[(f^{BMA}(\boldsymbol{x}))^{2}]-\mathbb{E}[(f^{BMA}(\boldsymbol{x}))]^{2}$.
II-B Adversarial Attacks
In a security-sensitive application, another priority, together with reliable uncertainty estimates, is to guarantee a certain level of robustness in a model that can be the target of attackers aiming at disrupting its normal functioning. This necessity brought to light what is known as adversarial machine learning, an ML field that keeps exploring the attack surfaces of an ML system to propose new countermeasures to achieve robustness [11]. In this regard, most of the research community efforts focused on attacks jeopardizing the integrity of a model, where an attacker is interested in changing a model’s prediction at test time [10, 9]. In such an attack scenario, the goal of the attacker is to craft a small perturbation $\boldsymbol{\delta}^{\prime}$ such that, once added to an input sample $\boldsymbol{x}$, it induces the target model $f^{\boldsymbol{\theta}}$ to output a wrong prediction, i.e., $f^{\boldsymbol{\theta}}(\boldsymbol{x}+\boldsymbol{\delta}^{\prime})\neq f^{\boldsymbol{\theta}}(\boldsymbol{x})$. Here, the perturbation is typically constrained to a limited domain $\mathcal{B}_{\epsilon}=\{\forall\boldsymbol{\delta}\ :||\boldsymbol{\delta}||_{p}\leq\epsilon\}$, where ||.$||.||_{p}$ is an arbitrary $L_{p}$ norm (e.g., $L_{1}$, $L_{2}$ or $L_{\infty}$) and $\epsilon$ is the perturbation budget indicating the limits of the attacker’s power.
@@ II-B Adversarial Attacks
Defending against adversarial attacks – Currently, the most promising approach to protect a model against adversarial attacks is to proactively anticipate the attacker by including them during training. Such a procedure, which takes the name of adversarial training (AT) [12], aims at finding the optimal parametrization $\boldsymbol{\theta}^{\prime}$ that ideally protects from all possible perturbations $\boldsymbol{\delta}\in\mathcal{B}_{\epsilon}$. This strategy can be formalized as the following minimax problem:
@@
where $\mathcal{L}$ is the objective that minimize classification error (e.g., the cross-entropy loss). Here, Eq. 6 is composed of (i) an inner maximization problem to find, for each sample, the perturbation $\boldsymbol{\delta}$ that achieves high loss, and (ii) an outer minimization problem to find the parameters $\boldsymbol{\theta}$ that minimize this worst-case loss.
III Adversarial Attacks Against Uncertainty Quantification
From the perspective of the most typical test-time evasion attacks, adversarial machine learning characterizes an attacker interested in making the model misclassify the prediction on the input adversarial example. In the context of Uncertainty Quantification, instead, the focus shifts from “deteriorating predictions” to “deteriorating the uncertainty measure”. The correct functioning of an uncertainty measure implies that the most likely correct predictions have a low uncertainty, which, in contrast, becomes high when the predictions are plausibly incorrect. Therefore, while predictions and uncertainty represent two different measures, they are, however, affine: the uncertainty estimate measures the likelihood that the given prediction is wrong. As we will see in the following subsection, state-of-the-art work conceived uncertainty adversarial attacks specifically to spoil this statistical correlation.
III-A Previous Work
@@
In our previous work [7], we proposed an alternative scenario in which the attacker is interested in increasing or decreasing the predictive uncertainty of all the input samples. The ability to damage predictive uncertainty without having specific information about the labels is an attractive feature, especially if such a measure is the input of a downstream module or is given to a human. For instance, imagine a system employing the uncertainty measure of an ML model for determining if the clinical data of a patient is sufficient to make a diagnosis: without knowledge of this specific domain (thus without the labels), an attacker may craft a perturbation for increase indiscriminately the uncertainty associated to the patient’s data. Such an attack may overload the processing pipeline of a hospital, causing the system to require more analysis, even when unnecessary. The proposed attack scenario takes the name of over-confidence (when decreasing the uncertainty) or under-confidence attack (when increasing the uncertainty). While for the over-confidence attack, we kept the same optimization process of  Eq. 8, the under-confidence case has been investigated only theoretically, without an actual loss function implementation.
@@
Obadinma et al.[8] propose to harm the model’s calibration, disrupting the functioning of any downstream module or human using the uncertainty measure. For doing so, they present four different attacks, two of which coincide with the formulation we proposed in  [7], i.e., (i) over-confidence and (ii) under-confidence attack. They further propose (iii) maximum miscalibration attack (MMA), where they increase or decrease the confidence of each sample following the same criteria of  [5], and (iv) random confidence attack (RCA), where they extend the latter but with a randomized confidence goal for each input, producing a more natural confidence distribution that, although being a less effective attack, results in being more stealthy. Based on these types of attacks against model’s calibration, they propose two ad-hoc defenses: (i) Calibration Attack Adversarial Training (CAAT), in which they perform adversarial training using samples generated with maximum miscalibration attacks, and (ii) Compression Scaling (CS), a post-processing technique where they heuristically distribute low scores to higher range. To the best of our knowledge, this is the only work that proposes a defense strategy against uncertainty attacks.
To the best of our knowledge, all state-of-the-art approaches rely exclusively on aleatoric uncertainty estimates derived from deterministic models, whereas epistemic uncertainty in the context of adversarial robustness has been explored solely in our prior work [7]. This common practice is not a deliberate choice, but rather a consequence of the unpopularity of Bayesian models among the adversarial training community, which makes robust models capable of providing principled epistemic measures hard to find.
III-B A Modular Unified Framework for Uncertainty Attacks
@@
Interestingly, all the state-of-the-art works build an optimization process with a proper loss formulation for accomplishing an attacker’s goal. This means that proving the model’s robustness against a specific goal does not directly translate to the same robustness guarantees on other goals. For example, analyzing the effect of an attack on OOD uncertainty minimization does not provide any information on the effectiveness against indiscriminate uncertainty maximization, making it difficult to compare models in terms of uncertainty robustness. Accordingly, we propose a modular approach for unifying the diverse attacks proposed in the literature, which, as a result, also unifies the notion of robustness. In general, all the possible instances of the attacker’s goal have a shared characteristic: they all need to modify the uncertainty measure associated with the model’s predictions; more sensitive uncertainty variations generally lead to less robustness against uncertainty attacks. As a result, any possible uncertainty attack discussed in literature (e.g., Maximum Miscalibration [8], Random Confidence [8], and OOD camoufladge [5]) can be traced back to a combination of over- and under-confidence attacks. Let us take random confidence as an example: knowing which is the maximum and minimum attainable uncertainty value under attack for a specific set of samples directly influences the success rate of the random confidence attack (the higher the possible excursion, the higher the space of action of the attack). As another example, given that for succeeding in a maximum miscalibration attack it is necessary to selectively increase or decrease the uncertainty of different samples, the attack success rate depends upon the robustness of the models against under- and over-confidence attacks.
Therefore, instead of focusing on specific instances of the attacker’s goal, we estimate the range of uncertainty values an attacker can reach when perturbing each input $x$ on the dataset $\mathcal{D}$. This approach provides a comprehensive overview of the model’s vulnerability under uncertainty manipulations, encompassing all possible attack scenarios. Concretely, given an attacker capable of perturbing a sample $\boldsymbol{x}$ inside a domain $\mathcal{B}_{\epsilon}$, we can distinguish between two different cases: (i) decreasing the uncertainty (i.e., over-confidence attack) and (ii) increasing the uncertainty (i.e., under-confidence attack). These objectives can be achieved, respectively, by minimising or maximising the uncertainty measure $\mathcal{U}(f^{\boldsymbol{\theta}}(\boldsymbol{x}))$:
@@
$\underbrace{\boldsymbol{\delta}^{\downarrow}=\operatornamewithlimits{\arg\,\max}_{\boldsymbol{\delta}\in\mathcal{B}_{\epsilon}}\;\mathcal{U}(f^{\boldsymbol{\theta}}(\boldsymbol{x}+\boldsymbol{\delta}))}_{\textbf{Over-confidence Attack}}\quad\underbrace{\boldsymbol{\delta}^{\uparrow}=\operatornamewithlimits{\arg\,\min}_{\boldsymbol{\delta}\in\mathcal{B}_{\epsilon}}\;\mathcal{U}(f^{\boldsymbol{\theta}}(\boldsymbol{x}+\boldsymbol{\delta}))}_{\textbf{Under-confidence Attack}}$ (9)
@@
In accordance with the state of the art, we use the entropy of the prediction vector $H(f^{\boldsymbol{\theta}}(\boldsymbol{x}))$ [1] as our uncertainty measure $\mathcal{U}(\cdot)$. To instantiate an over-confidence attack, the objective is to minimize the entropy of the output probability. To do so, it is sufficient to force the model to output $f^{\boldsymbol{\theta}}_{k}(\boldsymbol{x}+\boldsymbol{\delta}^{\downarrow})=1$ for an arbitrary $k\in\mathcal{Y}$. We know from [7] that the most straightforward strategy for the attacker is to maximize the probability of the class corresponding to the predicted class, as it already has the highest probability among all the classes, as in  Eq. 8; this attack takes the name of Stabilizing Attack (STAB) since it stabilizes the model’s prediction by enforcing the most likely class [7]. To achieve this objective, one can employ the cross-entropy loss, setting as target $\boldsymbol{t}^{\downarrow}$ the one-hot-encoding of the predicted label $\hat{y}$:
@@
IV Theoretical Analysis on Adversarial Uncertainty Robustness
Despite much literature on the attacker side, only one work dealt with adversarial robustness against uncertainty attacks, proposing two ad hoc strategies. In contrast to them, driven by the flourishing literature on defenses against prediction attacks (i.e., traditional adversarial examples), before thinking of ad hoc strategies, we ask ourselves if and to what extent such defenses, particularly adversarial training, may protect even against uncertainty attacks. Accordingly, our main research question is: